# Correlation functions of harmonic lattices in d-dimensional space

### v1

@@ Abstract
We study the correlation functions between the dynamical variables and between their conjugate momenta at sites of a harmonic lattice in the $d$-dimensional Euclidean space. We show that at the thermodynamic limit, they can be expressed in terms of Lauricella’s C-type hypergeometric series. Furthermore, using these expressions, we explicitly demonstrate that the correlators near the center of the lattice satisfying Diriclet boundary conditions coincide with those for the lattice with the periodic boundary conditions. By utilizing these expressions, we expect to make it easier to create programs that compute fast and highly precise for the quantum information quantities of subsystems within lattices.
@@ 1 Introduction
Indices of information loss due to entanglements, such as Renyi entropy are subjects of research not only in quantum information theory but also in many other fields, including field theory and condensed matter physics[1]. In the harmonic lattice model representing the discretized Klein Gordon field, to determine these quantities, one must calculate the vacuum correlators $Q_{\vec{r}\vec{s}}=\braket{q_{\vec{r}}q_{\vec{s}}}$ and $P_{\vec{r}\vec{s}}=\braket{p_{\vec{r}}p_{\vec{s}}}$ for the set of sites of interest, where $q_{\vec{r}},q_{{\vec{s}}}$ and $p_{\vec{r}},p_{\vec{s}}$ are quantized dynamical variables of the sites with position vectors $\vec{r},\vec{s}$ and their conjugate momentum operators, respectively. Hereafter, we will simply refer to these correlators as $Q$ and $P$ or matrices $Q$,$P$ and so on as needed.
For example, consider an isolated lattice $U$ and a subset $A$ of sites, where we wish to obtain Renyi entropy of degree $n$, $S_{A}^{(n)}$, between A and its complement. The procedure is as follows: First, compute the matrices $Q_{\vec{r}\vec{s}}$ and $P_{\vec{r}\vec{s}}$ restricted to sites $\vec{r}$ and $\vec{s}$ within $A$. Second, find the eigenvalues $\xi_{i}^{2}$ of $\Xi=QP$. The Renyi entropy can be computed using the following formula[2].
@@
In this manuscript, we show that by considering the thermodynamic limit, the lattice scale $L\to\infty$ and replacing the sum over all site positions $(k_{1},k_{2},...,k_{d})$ within $U$ appearing in the computation of $Q$ and $P$ with multiple integrals $\int\int\dots\int dk_{1}dk_{2}\dots dk_{d}$ and calculating them analytically, the results for spatial dimensions of two or more are expressed by the C-type Lauricella’s hypergeometric series, $F_{C}$, irrespective of the boundary conditions for the dynamical variables $q_{r}$s and their conjugate momenta $p_{r}$s. We expect that by utilizing these expressions along with appropriate numerical computation libraries, it makes easier to create programs that rapidly compute various information quantities arising from entanglement.
@@ 2 Correlators on finite harmonic lattices
We consider a free scalar field in $(d+1)$ dimensional Lorentzian space-time and discretize it into a cubic lattice with $L^{d}$ sites of harmonic oscillators coupled with neighbor sites by spring-like interaction. The Hamiltonian is given by
$H=\frac{1}{2}\sum_{\vec{j}}(p_{\vec{j}}^{2}+\omega^{2}q_{\vec{j}}^{2}+\sum_{\alpha=1}^{d}(q_{\vec{j}+\vec{e}_{\alpha}}-q_{\vec{j}})^{2}),$ (2)
where $\vec{j}=(j_{1},j_{2},\ldots,j_{d})$ denotes the position of a site and $\vec{e}_{\alpha}=(e_{1}=0,0,\dots,e_{\alpha}=1,0,\ldots,0)$ is a unit vector on a spatial axis.
When the lattice $U$ satisfies periodic boundary conditions(PBCs) along directions of spatial axes, $q_{\vec{j}+L\vec{e}_{\alpha}}=q_{\vec{j}},\ p_{\vec{j}+L\vec{e}_{\alpha}}=p_{\vec{j}},\ \alpha=1,2,\ldots,d$, based on the variables $q_{\vec{j}}$s and $p_{\vec{j}}$s obtained by solving the Hamiltonian equations, the dispersion relation $\omega_{k}$ and $Q$, $P$ matrices are given [4] by
$\omega_{\vec{k}}^{(P)}$ $=\sqrt{\omega^{2}+4\sum_{\alpha=1}^{d}\sin^{2}(\pi k_{\alpha}/L)}\ ,$ (3)
$Q_{\vec{r}\vec{s}}=\braket{q_{\vec{r}}q_{\vec{s}}}$ $=\frac{1}{2V}\sum_{\vec{k}}\frac{1}{\omega_{\vec{k}}^{(P)}}\prod_{\alpha=1}^{d}\cos[2\pi k_{\alpha}(r_{\alpha}-s_{\alpha})/L],$ (4)
$P_{\vec{r}\vec{s}}=\braket{p_{\vec{r}}p_{\vec{s}}}$ $=\frac{1}{2V}\sum_{\vec{k}}\omega_{\vec{k}}^{(P)}\prod_{\alpha=1}^{d}\cos[2\pi k_{\alpha}(r_{\alpha}-s_{\alpha})/L],$ (5)
where $k_{\alpha}$’s are integers such that $0\leq k_{\alpha}\leq L-1$ and $V=L^{d}$.
@@
$\omega_{\vec{k}}^{(D)}$ $=\sqrt{\omega^{2}+4\sum_{\alpha=1}^{d}\sin^{2}(\pi k_{\alpha}/(2L))},$ (6)
@@ 3 Correlation functions in the thermodynamic limit
$G^{(\pm 1/2)}_{\vec{r}\vec{s}}=\frac{\Omega^{\mp 1}}{2(2\pi)^{d}}\sum_{m=0}^{\infty}\frac{(\pm 1/2)_{m}}{m!}v^{m}\underbrace{\int_{0}^{2\pi}\dots\int_{0}^{2\pi}d^{d}k\ \left(\sum_{\alpha=1}^{d}\cos k_{\alpha}\right)^{m}\prod_{\alpha^{\prime}=1}^{d}\cos n_{\alpha^{\prime}}k_{\alpha^{\prime}}}_{I_{m,n_{1},n_{2},\ldots,n_{d}}},$ (11)
The multiple integral $I_{m,n_{1},n_{2},\ldots,n_{d}}$ in (11) can be expressed as a linear combination of the products of integrals with respect to each $k_{\alpha}$ by expanding $(\sum_{\alpha=1}^{d}\cos k_{\alpha})^{m}$ as follows:
@@
$I_{l_{\alpha},n_{\alpha}}=\int_{0}^{2\pi}\cos^{l_{\alpha}}k_{\alpha}\cos n_{\alpha}k_{\alpha}dk_{\alpha}=\begin{cases}\frac{2\pi}{2^{l_{\alpha}}}\begin{pmatrix}l_{\alpha}\\ (l_{\alpha}-n_{\alpha})/2\end{pmatrix}&\text{if}\ l_{\alpha}\geq n_{\alpha}\ \text{and}\ l_{\alpha}=n_{\alpha}\hskip-5.69054pt\mod 2,\\ 0&\text{otherwise}.\end{cases}$ (13)
@@
The infinite series appearing in equation (16) are Lauricella’s C-type hypergeometric series defined by:
@@
The difference between (9) and (24), besides $n_{\alpha}$ changing to $n_{\epsilon_{\alpha}}$, are that (i)the coefficient before the integral changes from $1/(2\cdot 2^{d}\pi^{d})$ to $1/(2\pi^{d})$ and that (ii) the integration range for each $k_{\alpha}$ changes from $0,2\pi]$ to $0,\pi]$. Among these, due to (ii), the value of the integral for each $k_{\alpha}$ in (24) becomes half that of (13), and the entire multiple integral becomes multiplied by $1/2^{d}$. This cancels exactly with the change due to (i), so (24) matches (9) and its calculated result (16) except for the substitution from $n_{\alpha}$ to $n_{\epsilon_{\alpha}}$:
@@
where we set $N\overset{\underset{\mathrm{def}}{}}{=}n_{\epsilon_{1}}+n_{\epsilon_{2}}+\dots+n_{\epsilon_{d}}$, $u\overset{\underset{\mathrm{def}}{}}{=}p_{1}+p_{2}+\dots p_{d}$ as before. Then correlation functions $Q$ and $P$ under DBCs can be expressed, respectively, using Lauricella’s hypergeometric series as follows:
@@
Thus, we have shown that for harmonic lattices in space with $d\geq 2$, the correlation functions for the dynamical variables in the thermodynamic limit are expressed as a hypergeometric series of the same type, regardless of the boundary conditions of the lattice.
@@ 4 Connection between correlators of different boundary conditions
$(a)_{N}$ $\sim(N_{+}+N_{-}+a)^{N_{+}+N_{-}+a-1/2}e^{-(N_{+}+N_{-}+a)}\sim N_{+}^{N_{+}}\left(1+\frac{N_{-}+a}{N_{+}}\right)^{N_{+}}N_{+}^{N_{-}+a-1/2}e^{-N_{+}}$
$\sim N_{+}^{N_{+}}e^{N_{-}}e^{-N_{+}}N_{+}^{N_{-}+a-1/2}.$ (28)
@@
$(\prod_{\alpha=1}^{d}n_{\alpha}!)^{-1}\sim(\prod_{\alpha=1}^{d_{1}}n_{\alpha}^{n_{\alpha}+1/2}e^{-N_{+}})^{-1}\sim(\prod_{\alpha=1}^{d_{1}}\lambda_{\alpha}^{\lambda_{\alpha}N_{+}})^{-1}N_{+}^{-N_{+}}N_{+}^{-d_{1}/2}e^{N_{+}}.$ (29)
@@
$(\prod_{\alpha=1}^{d}n_{\alpha}!)^{-1}\sim d_{1}^{N_{+}}N_{+}^{-N_{+}}N_{+}^{-d_{1}/2}e^{N_{+}}.$ (30)
@@
Since each $(p_{1},p_{2},\dots,p_{d})$ term of infinite series in (25) is obtained in the thermodynamic limit ($n^{(\alpha)}_{1}\sim L\to\infty$), $p_{\alpha}$s in the term should be regarded as finite values. So the absolute value of (33) decreases exponentially and becomes zero. Thus, only $J_{n_{0}n_{0},\ldots,n_{0}}$ survives on the right-hand side of (23), and the correlation functions for the DBC case $J_{\vec{r}\vec{s}}^{(\pm 1/2)}$ coincide with that for the PBC case, namely $G_{\vec{r}\vec{s}}^{(\pm 1/2)}$ in (16).
@@ 5 Conclusions and future tasks
For harmonic lattices, we have shown that the vacuum correlation function matrices $Q$ and $P$ for the dynamical variables and their conjugate momenta on sites are consistently expressed using Lauricella’s C-type hypergeometric series $F_{C}^{(d)}$ in the thermodynamic limit, in any spatial dimension $d$ of the lattice. For lattices satisfying periodic boundary conditions, $Q_{\vec{r}\vec{s}}$ and $P_{\vec{r}\vec{s}}$ are given by (18) and (19), and for those satisfying the Dirichlet boundary conditions, they are given by (26) and (27). These hypergeometric series absolutely converge for spatial dimensions $d\geq 2$. Additionally, we have specifically observed that the correlation functions $Q_{\vec{r}\vec{s}}$ and $P_{\vec{r}\vec{s}}$ for sites near the center of the lattice under Dirichlet boundary conditions coincide with those in the lattice under periodic boundary conditions. In the specific cases, it is known that Lauricella’s C-type hypergeometric series is factorized into the product of Gauss’s hypergeometric series $_{2}F_{1}$s. As an example, let us consider $F_{C}^{(2)}$ appearing in the $Q$ of (18) for $d=2$ which is also also referred as Appell’s $F_{4}$ hypergeometric series. For the massless harmonic lattice, the variables of this hypergeometric series are $v^{2}=1/4$ and it is factorized as follows[5]:
@@
Numerical calculations of $Q$ and $P$ matrices are essential for research using harmonic lattices on entanglement indices like Renyi entropy and their corner functions, edge terms and contour functions, which represents the contribution to the indices from vertices, edges on the boundary of the physical system and from each site within it. However, as the dimensionality increases, creating programs to perform these computations quickly under memory constraints for sufficiently large-scale subsystems becomes increasingly difficult. The analytical solutions presented here, which express the correlation functions in $d$-dimensional space using hypergeometric series, are expected to facilitate the creation of these programs.
The following two issues remain: 1.We will investigate whether (34) and other hypergeometric series in correlator functions can be written in more concise forms. 2.For $d=2$ and $d=3$, we will obtain corner functions and edge terms for specific shapes and scales of subsystems based on the $Q$ and $P$ values computed by the program cord that employs expressions (18) and (19). Then, we will compare these results with those from previous studies, which were computed based on the $Q$ and $P$ values in the finite-scale lattice of (4) and (5) or on those of integral expressions in (11).

### v2

@@ Abstract
We study the correlation functions between the dynamical variables and between their conjugate momenta at sites of a harmonic lattice in the $d$-dimensional Euclidean space. We show that at the thermodynamic limit, they can be expressed in terms of Lauricella’s C-type hypergeometric functions. Furthermore, using these expressions, we explicitly demonstrate that the correlators near the center of the lattice satisfying Diriclet boundary conditions coincide with those for the lattice with the periodic boundary conditions. By utilizing these expressions, we expect to make it easier to create programs that compute fast and highly precise for the quantum information quantities of subsystems within lattices.
@@ 1 Introduction
Indices of information loss due to entanglements, such as Rényi entropy are subjects of research not only in quantum information theory but also in many other fields, including field theory and condensed matter physics[1]. In the harmonic lattice model representing the discretized Klein Gordon field, these indices are determined from the vacuum correlators $Q_{\vec{r}\vec{s}}=\braket{q_{\vec{r}}q_{\vec{s}}}$ and $P_{\vec{r}\vec{s}}=\braket{p_{\vec{r}}p_{\vec{s}}}$ for the set of sites of interest, where $q_{\vec{r}},q_{{\vec{s}}}$ and $p_{\vec{r}},p_{\vec{s}}$ are quantized dynamical variables of the sites with position vectors $\vec{r},\vec{s}$ and their conjugate momentum operators, respectively. Hereafter, we will simply refer to these correlators as $Q$ and $P$ or $Q$,$P$ matrices as needed.
For example, consider an isolated lattice $U$ and its subset $A$ of sites, where we evaluate Rényi entropy $S_{A}^{(n)}$, between $A$ and its complement. The procedure is as follows: First, compute the matrices $Q_{\vec{r}\vec{s}}$ and $P_{\vec{r}\vec{s}}$ restricted to sites $\vec{r}$ and $\vec{s}$ belonging to $A$. Second, find the eigenvalues $\xi_{i}^{2}$ of $\Xi=QP$. The Rényi entropy of degree $n$ can be computed using the following formula[2].
@@
In this manuscript, we show that by considering the thermodynamic limit, the lattice scale $L\to\infty$ and replacing the sum over all site positions $(k_{1},k_{2},\ldots,k_{d})$ within $U$ appearing in the computation of $Q$ and $P$ with multiple integrals $\int\int\dots\int dk_{1}dk_{2}\dots dk_{d}$ and calculating them analytically, the results for any spatial dimension expressed by the C-type Lauricella’s hypergeometric functions, $F_{C}$, irrespective of the boundary conditions for the dynamical variables $q_{\vec{r}}$s and their conjugate momenta $p_{\vec{r}}$s. We expect that by utilizing these expressions along with appropriate numerical computation libraries, it makes easier to create programs that rapidly compute various information quantities arising from entanglement.
@@ 2 Correlators on finite harmonic lattices
We consider a free scalar field in $(d+1)$ dimensional Lorentzian space-time and discretize it into a cubic lattice with $L^{d}$ sites of harmonic oscillators coupled with nearest neighborhood. The Hamiltonian is given by
$H=\frac{1}{2}\sum_{\vec{j}}\left(p_{\vec{j}}^{2}+\omega^{2}q_{\vec{j}}^{2}+\sum_{\alpha=1}^{d}(q_{\vec{j}+\vec{e}_{\alpha}}-q_{\vec{j}})^{2}\right),$ (2)
where $\vec{j}=(j_{1},j_{2},\ldots,j_{d})$ denotes the position of a site and $\vec{e}_{\alpha}=(0,0,\ldots,0,1,0,\ldots,0)$ is a unit vector in the $\alpha$ direction.
When the lattice $U$ satisfies periodic boundary conditions(PBCs) along directions of spatial axes, $q_{\vec{j}+L\vec{e}_{\alpha}}=q_{\vec{j}},\ p_{\vec{j}+L\vec{e}_{\alpha}}=p_{\vec{j}},\ \alpha=1,2,\ldots,d$, solving Hamilton’s canonical equations yields the variables $q_{\vec{j}}$s and $p_{\vec{j}}$s. Quantizing these then, the dispersion relation $\omega_{k}$ and $Q$, $P$ matrices are given [4] by
$\omega_{\vec{k}}^{(P)}$ $=\sqrt{\omega^{2}+4\sum_{\alpha=1}^{d}\sin^{2}\frac{\pi k_{\alpha}}{L}}\ ,$ (3)
$Q_{\vec{r}\vec{s}}=\braket{q_{\vec{r}}q_{\vec{s}}}$ $=\frac{1}{2V}\sum_{\vec{k}}\frac{1}{\omega_{\vec{k}}^{(P)}}\prod_{\alpha=1}^{d}\cos\frac{2\pi k_{\alpha}(r_{\alpha}-s_{\alpha})}{L},$ (4)
$P_{\vec{r}\vec{s}}=\braket{p_{\vec{r}}p_{\vec{s}}}$ $=\frac{1}{2V}\sum_{\vec{k}}\omega_{\vec{k}}^{(P)}\prod_{\alpha=1}^{d}\cos\frac{2\pi k_{\alpha}(r_{\alpha}-s_{\alpha})}{L},$ (5)
where $k_{\alpha}$s are integers such that $0\leq k_{\alpha}\leq L-1$ and $V=L^{d}$.
@@
$\omega_{\vec{k}}^{(D)}$ $=\sqrt{\omega^{2}+4\sum_{\alpha=1}^{d}\sin^{2}\frac{\pi k_{\alpha}}{2L}},$ (6)
@@ 3 Correlation functions in the thermodynamic limit
$G^{(\pm 1/2)}_{\vec{r}\vec{s}}=\frac{\Omega^{\mp 1}}{2(2\pi)^{d}}\sum_{m=0}^{\infty}\frac{(\pm 1/2)_{m}}{m!}v^{m}\int_{0}^{2\pi}\dots\int_{0}^{2\pi}d^{d}k\ \left(\sum_{\alpha=1}^{d}\cos k_{\alpha}\right)^{m}\prod_{\alpha^{\prime}=1}^{d}\cos n_{\alpha^{\prime}}k_{\alpha^{\prime}}.$ (11)
Here, we denote the multiple integral appearing in (11) as $I_{m,n_{1},n_{2},\ldots,n_{d}}$. Then it can be expressed as a linear combination of the products of integrals with respect to each $k_{\alpha}$ by expanding $(\sum_{\alpha=1}^{d}\cos k_{\alpha})^{m}$ as follows:
@@
$I_{l_{\alpha},n_{\alpha}}=\int_{0}^{2\pi}\cos^{l_{\alpha}}k_{\alpha}\cos n_{\alpha}k_{\alpha}dk_{\alpha}=\begin{cases}\displaystyle\frac{2\pi}{2^{l_{\alpha}}}\begin{pmatrix}l_{\alpha}\\ (l_{\alpha}-n_{\alpha})/2\end{pmatrix}&\text{if}\ l_{\alpha}\geq n_{\alpha}\ \text{and}\ l_{\alpha}=n_{\alpha}\hskip-5.69054pt\mod 2,\\ 0&\text{otherwise}.\end{cases}$ (13)
@@
The infinite series appearing in equation (16) is Lauricella’s C-type hypergeometric series defined by:
@@
The differences between (9) and (24), besides $n_{\alpha}$ changing to $n_{\epsilon_{\alpha}}$, are that (i) the coefficient before the integral changes from $1/(2\cdot 2^{d}\pi^{d})$ to $1/(2\pi^{d})$ and that (ii) the integration range for each $k_{\alpha}$ changes from $0,2\pi]$ to $0,\pi]$. Among these, due to (ii), the value of the integral for each $k_{\alpha}$ in (24) becomes half that of (13), and the entire multiple integral becomes multiplied by $1/2^{d}$. This cancels exactly with the change due to (i), so (24) matches (9) except for the substitution of $n_{\epsilon_{\alpha}}$ for $n_{\alpha}$. Thus, Eq.(24) leads to
@@
where we set $N\overset{\underset{\mathrm{def}}{}}{=}n_{\epsilon_{1}}+n_{\epsilon_{2}}+\dots+n_{\epsilon_{d}}$, $u\overset{\underset{\mathrm{def}}{}}{=}p_{1}+p_{2}+\dots p_{d}$ as before. Then correlation functions $Q$ and $P$ under DBCs can be expressed, respectively, using Lauricella’s hypergeometric functions as follows:
@@
Thus, we have shown that for harmonic lattices in space with $d\geq 2$, the correlation functions in the thermodynamic limit are expressed by hypergeometric functions of the same type, regardless of the boundary conditions of the lattice.
@@ 4 Connection between correlators of different boundary conditions
$(a)_{N}$ $\sim(N_{+}+N_{-}+a)^{N_{+}+N_{-}+a-1/2}e^{-(N_{+}+N_{-}+a)}\sim N_{+}^{N_{+}}\left(1+\frac{N_{-}+a}{N_{+}}\right)^{N_{+}}N_{+}^{N_{-}+a-1/2}e^{-N_{+}-N_{-}}$
$\sim N_{+}^{N_{+}}e^{-N_{+}}N_{+}^{N_{-}+a-1/2}.$ (28)
@@
$(\prod_{\alpha=1}^{d}n_{\alpha}!)^{-1}\sim(e^{-N_{+}}\prod_{\alpha=1}^{d_{1}}n_{\alpha}^{n_{\alpha}+1/2}e^{-N_{+}})^{-1}\sim(\prod_{\alpha=1}^{d_{1}}\lambda_{\alpha}^{\lambda_{\alpha}N_{+}})^{-1}N_{+}^{-N_{+}}N_{+}^{-d_{1}/2}e^{N_{+}}.$ (29)
@@
$(\prod_{\alpha=1}^{d}n_{\alpha}!)^{-1}\leq d_{1}^{N_{+}}N_{+}^{-N_{+}}N_{+}^{-d_{1}/2}e^{N_{+}}.$ (30)
@@
Since each $(p_{1},p_{2},\dots,p_{d})$ term of infinite series in (25) is obtained in the thermodynamic limit ($n^{(\alpha)}_{1}\sim L\to\infty$), $p_{\alpha}$s in the term should be regarded as finite values. So the absolute value of (33) decreases exponentially and becomes zero. Thus, only $J_{n_{0},n_{0},\ldots,n_{0}}$ survives on the right-hand side of (23), and the correlation functions for the DBC case $J_{\vec{r}\vec{s}}^{(\pm 1/2)}$ coincide with that for the PBC case, namely $G_{\vec{r}\vec{s}}^{(\pm 1/2)}$ in (16).
@@ 5 Conclusions and future tasks
For harmonic lattices, we have shown that the vacuum correlation function matrices $Q$ and $P$ for the dynamical variables and their conjugate momenta on sites are consistently expressed using Lauricella’s C-type hypergeometric series $F_{C}^{(d)}$ in the thermodynamic limit, in any spatial dimension $d$ of the lattice. For lattices satisfying periodic boundary conditions, $Q_{\vec{r}\vec{s}}$ and $P_{\vec{r}\vec{s}}$ are given by (18) and (19), and for those satisfying the Dirichlet boundary conditions, they are given by (26) and (27). These hypergeometric series absolutely converge for spatial dimensions $d\geq 2$. Additionally, we have specifically observed that the correlation functions $Q_{\vec{r}\vec{s}}$ and $P_{\vec{r}\vec{s}}$ for sites near the center of the lattice under Dirichlet boundary conditions coincide with those in the lattice under periodic boundary conditions.
In the specific cases, it is known that Lauricella’s C-type hypergeometric functions is factorized into the product of Gauss’s hypergeometric function $_{2}F_{1}$s. As an example, let us consider $F_{C}^{(2)}$ appearing in the $Q$ of (18) for $d=2$ which is also referred as Appell’s $F_{4}$ hypergeometric functions. For the massless harmonic lattice, the arguments of this hypergeometric functions are $v^{2}=1/4$ and it is factorized as follows[5]:
@@
Numerical calculations of $Q$ and $P$ matrices are essential for research using harmonic lattices on entanglement indices like Rényi entropy and their corner functions, edge terms and contour functions, which represents the contribution to the indices from vertices, edges on the boundary of the physical system and from each site within it. However, as the dimensionality increases, creating programs to perform these computations quickly under memory constraints for sufficiently large-scale subsystems becomes increasingly difficult. The analytical solutions presented here, which express the correlation functions in $d$-dimensional space using hypergeometric functions, are expected to facilitate the creation of these programs.
The following two issues remain: 1.We will investigate whether hypergeometric functions in correlator functions in this manuscript can be written in more concise forms as the factorization (34). 2.For $d=2$ and $d=3$, we will obtain corner functions and edge terms for specific shapes and scales of subsystems based on the $Q$ and $P$ values computed by the program cord that employs expressions (18) and (19). Then, we will compare these results with those from previous studies, which were computed based on the $Q$ and $P$ values in the finite-scale lattice of (4) and (5) or on those of integral expressions in (11).